OpenStudy (anonymous):

expand sin (x-x^2) in power series until term containing x^4

OpenStudy (anonymous):

The quick way would be to use the power series for \(\sin x\): \[\sin x=\sum_{n=0}^\infty \frac{(-1)^nx^{2n+1}}{(2n+1)!}\] Substituting \(x-x^2\) for \(x\), you have \[\sin (x-x^2)=\sum_{n=0}^\infty \frac{(-1)^n(x-x^2)^{2n+1}}{(2n+1)!}\] Then plug in as many \(n\) as you need to get a term containing \(x^4\).

OpenStudy (anonymous):

Thanks! Btw, can I express this by using limits?

OpenStudy (reemii):

There is a little problem with the answer above. different terms (for different n's) will contain "x" terms of same degree.

OpenStudy (reemii):

hm. I'm wrong actually, but the power series must be of the form \(\sum a_k x^k\). each \(k\) deals with only one \(x^k\).

OpenStudy (anonymous):

can i express in the form of maclaurin series?

OpenStudy (reemii):

You need either to compute completely all derivatives of the function you are given, or use another "better trick". use that: \(\sin(y) = y-y^3/3! + y^5/5! -...\) and put \(y=x-x^2\). After rearranging terms, you will have it in the form \(sin(x-x^2) = ..x + ..x^2 + ...x^3 + ...x^4 + ...\).

OpenStudy (reemii):

Actually we only need those 2 terms: \(y-y^3/3!+ -..\), with y=x-x^2, because \(y^5\) will give you \(x^5\) and above, and you don't need that. here: \(y-y^3/3! = (x-x^2) - (x-x^2)^3/3! = x-x^2 - [x^3-3x^2x^2 + 3xx^4 - x^6]/3!\). you can forget the \(x^5\) and the \(x^6\) term.

OpenStudy (anonymous):

I'm a bit confuse. How to get this? \[\sin(y)=y−y3/3!+y5/5!−...\] Any name for this methods? Thanks :)

OpenStudy (reemii):

This is a "well known" formula. I would say that if you study Taylor studies, this is one you should remember, also the cos and the \(e^x\) function.

OpenStudy (reemii):

But you can obtain the series by using the taylor formula.. \(\sum_n \frac{f^{(n)}(x_0)}{n!}(x-x_0)^n\)

OpenStudy (reemii):

Taylor series*

OpenStudy (reemii):

and the \(\frac{1}{1-x}\) series.

OpenStudy (anonymous):

So all I need to do is to express \[y= (x-x^2)\] and sub into the series later? And can I directly use maclaurin series? since it's just a special case for taylor's series.

OpenStudy (reemii):

yes. this works well, but you need to be certain that you CANNOT obtain any terms x,x^2,x^3, or x^4 in the following terms. Here, we see easily that y^5 will NOT give any term lower than x^5, therefore we can forget this term and all the following ones.

OpenStudy (reemii):

maybe this is called substitution method. to be verified

OpenStudy (reemii):

there is an example of this technique on wiki https://en.wikipedia.org/wiki/Taylor_series#Calculation_of_Taylor_series

OpenStudy (anonymous):

I try directly differentiating the value \[\sin (x-x^2)\] but the value i get is 0 for x^2 and x^4. I think your way is the best solution so far.